# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v3

@@ VI Applications
This section explores literature that employs uncertainty-based downstream tasks on segmentation models. These include estimating the segmentation mask distribution subject to observer variability (Section VI-A), model introspection (i.e. ability to self-assess, Section VI-B), improved generalization (Section VI-C), and reduced labeling costs using Active Learning (Section VI-D).
VI-A Observer variability
@@
For evaluation, a commonly used metric minimizes the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick [129]. This metric is known as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) [130], denoted as
$\text{GED}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]$
$\hskip 10.0pt+\mathbb{E}_{\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}\sim P_{\hat{\mathbf{Y}}}}[k(\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime})]$
$\hskip 20.0pt-2\cdot\mathbb{E}_{\mathbf{y}\sim P_{\mathbf{Y}}}\mathbb{E}_{\hat{\mathbf{y}}\sim P_{\hat{\mathbf{Y}}}}[k(\mathbf{y},\hat{\mathbf{y}})],$ (37)
with marginals $P_{\mathbf{Y}}$ and $P_{\hat{\mathbf{Y}}}$, representing the true and predictive segmentation distribution, and a distance represented by kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the (1--IoU) or (1--Dice) score (squared by the GED). An alternative metric known as Hungarian Matching (HM) compares the predictions against the ground-truth labels through a cost matrix [74]. Subsequently, the unique optimal coupling between the two sets that minimize the average cost is determined through a combinatorial optimization algorithm. This can be also formally denoted as finding the permutation matrix $\mathbf{P}$, subject to the objective

### v4

@@ 6 Applications
This section briefly provides literature that incorporates uncertainty into downstream tasks using segmentation models. These tasks can be broadly categorized into four main groups: estimating the segmentation mask distribution subject to observer variability (Section 6.1), reducing labeling costs using Active Learning (Section 6.2), model introspection (i.e. ability to self-assess Section 6.3) and improved generalization (Section 6.4). For quick reference, Table 5 summarizes the types of dataset domains (e.g., medical, outdoor scenes) associated with various downstream tasks.
6.1 Observer variability
@@
Evaluation
For evaluation, a commonly used metric minimizes the squared distance between arbitrary mean embeddings of the ground truth and predicted annotations using the kernel trick (Shawe-Taylor, 2004). This metric is known as Maximum Mean Discrepancy (MMD) or the Generalized Energy Distance (GED) (Gretton et al., 2012), denoted as
$\text{GED}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]+\mathbb{E}_{\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}\sim P_{\hat{\mathbf{Y}}}}[k(\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime})]-2\cdot\mathbb{E}_{\mathbf{y}\sim P_{\mathbf{Y}}}\mathbb{E}_{\hat{\mathbf{y}}\sim P_{\hat{\mathbf{Y}}}}[k(\mathbf{y},\hat{\mathbf{y}})],$ (36)
[…]